Please help
A Radio Tuning Circuit Part A of the L-C circut is 1.70 MHx, corresponding to one end of What is the inductance L of a coll connected to this capacitor i the oscillation frequency A rado can be tuned into a particular station frequency by adjusting the capacitance in an 1-C drcuit. Suppose that the minimum capacitance of a variable capackoe in a radio b 4 20 AM radio broadcast band, when the capacar is set to its minimum capidance? Express your answer in henrys to three significant figures My Answers Give Up The frequency at the other end of the broadcast band is 0 540 MEHe. What is the maximum capacitance C tequency stable over te range of te broadeast bard? of the capacifor if the oscillation Express your answer in farads to three significant figures
Answer #1

(A) XL = XC

2 pi f L = 1 / (2 pi f C)

4 x pi^2 x (1.70 x 10^6)^2 L x 4.20 x 10^-12 = 1

L = 2.09 x 10^-3 H .......Ans

(B) 4 x pi^2 x (0.540 x 10^6)^2 x 2.09 x 10^-3 x C = 1

C = 41.6 x 10^-12 F